# Enhancing Perception Quality in Remote Sensing Image Compression via Invertible Neural Network

### v1

@@ 3.1.1 Invertible block
Besides, as illustrated in Figure 2 (b), the forward network outputs the synthetic image $\mathbf{s}$ and the latent variable $\mathbf{z}$. Following the studies [26, 23], we assume that $\mathbf{z}$ follows a Gaussian distribution, which can be randomly sampled from the same distribution for the inverse mapping. Consequently, in the training stage, the synthetic image $\mathbf{s}$ and $\mathbf{z}$ can be merged and fed into the inverse network for model optimization. In the inference stage, the decoded image $\tilde{\mathbf{x}}$ and resampled $\vec{\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I})$ can be combined and fed into the inverse network for image enhancement.
3.1.2 Channel expansion and Haar transformation
Within each invertible block, the channels of both its input and output signals have the same number. However, in addition to generating the synthetic image, we also need to produce an expanded output to derive the Gaussian latent variable $\mathbf{z}$. To achieve this, as shown in Figure 2 (b), we double the number of channels of the input image by using $1\times 1$ convolutional layers. In addition, Haar transformation [29] is used here to separate the compression distortion information by splitting the high and low frequencies, which are then fed into the invertible blocks.
@@ 3.1.2 Channel expansion and Haar transformation
$\mathbf{p}_{i}=\operatorname{Split}(H_{D}^{-1}(\mathbf{F})),~{}~{}i\in\{1,2\};~{}~{}~{}~{}\mathbf{x}=\Phi^{-1}_{1}(\mathbf{p}_{1})+\Phi^{-1}_{2}(\mathbf{p}_{2}),~~$ (4)
@@
3.1.3 Conditional generation module (CGM)
As illustrated above, we aim to encode the compression distortion information of decoded RS images into a set of Gaussian-distributed latent variables, where the mean and variance are conditioned on the synthetic image. This conditioning enables the reconstruction process to be image-adaptive. To achieve this, as illustrated in Figure 2 (b), we introduce a CGM at the end of the forward network. Specifically, to establish the dependency between $\mathbf{s}$ and $\mathbf{z}$, as depicted in Figure 3, the output six-channel tensor $\mathbf{g}$ from the last invertible block is further divided by CGM into two parts: a three-channel synthetic image $\mathbf{s}$ and a three-channel latent variable $\tilde{\mathbf{z}}$ representing the compression distortion information.
@@ 3.1.3 Conditional generation module (CGM)
$\bm{\mu}=\psi_{1}(\mathbf{s});~{}~{}~{}~{}\bm{\sigma}=\exp(\psi_{2}(\mathbf{s})),$ (5)
@@
3.1.4 Quantization module (QM)
To ensure compatibility with common decoded image storage formats such as RGB (8 bits for each R, G, and B color channels), as shown in Figure 2 (b), we integrate QM after CGM. This module converts the floating-point values of the produced synthetic images into 8-bit unsigned integers by a rounding operation for quantization. However, it’s essential to acknowledge a significant obstacle: the quantization module is inherently non-differentiable. To address this challenge, we use the Straight-Through Estimator technique as employed in [15] to ensure that INN-RSIC is effectively optimized during the training process. Subsequently, in the inference stage, the decoded image can be reasonably fed into the inverse network for image enhancement.
3.2 Optimization strategy
3.2.1 Compression optimization loss
As we aim to capture the compression distortion of ELIC [15], here we provide a brief overview of its loss function. Specifically, to balance the compression ratio and the image quality of the decoded image of ELIC, the loss function is defined as
@@ 3.2.1 Compression optimization loss
3.2.2 Forward presentation loss
@@
Decoded image loss. To generate the guidance images for training the forward network, ELIC is adopted to generate the labeled image $\tilde{\mathbf{x}}$, which can be formulated as
@@
Case-agnostic distribution loss. To regularize the distribution of $\mathbf{z}$, we maximize the log-likelihood of $p(\mathbf{z})$, inspired by [25]. Thus, the loss function to constrain the latent variables $\mathbf{z}$ can be formulated as
$L_{\text{case}}=-\log(p(\mathbf{z}))~=-\log\left(\frac{1}{(2\pi)^{M/2}}\exp\left(-\frac{1}{2}||\mathbf{z}||^{2}\right)\right),$ (9)
where $M$ is the dimensionality of $\mathbf{z}$. This constraint penalizes the normalized latent variables $\mathbf{z}$ to follow standard Gaussian distribution. Consequently, in the inverse reconstruction processing, we can randomly sample a set of Gaussian-distributed variables along with the synthetic image to obtain the reconstructed image.
3.2.3 Reverse reconstruction loss
@@
Enhancing reconstruction loss. In theory, the synthetic image can be perfectly restored to the corresponding ground truth version through the inverse network of INN because there is no information omission. However, in practice, the decoded image is not generated by the forward network of our INN-RSIC but with an image compression method. That is, the synthetic image should be stored in 8-bit unsigned integer format so that the decoded image can be used directly instead of the synthetic image here for image detail enhancement during the inference stage. To address this, we adopt QM at the end of the forward network during training. Consequently, by penalizing the discrepancy between the reconstructed image and the ground truth, we can derive the enhancing reconstruction loss given as
@@
Quality perception loss. To improve the perceptual performance of the network by estimating distances in predefined feature space rather than image space, a perceptual loss function called $L_{\text{per}}$ is used here. In other words, the feature space distance is represented by an optimization function, which serves as a driver for the network to perform image reconstruction while retaining a feature representation similar to the ground truth. Concretely, the learning perceptual image patch similarity (LPIPS) [38] is utilized to indicate the perceptual loss, which is defined as
@@
where $\operatorname{Vgg}(\cdot)$ denotes a feature function that leverages the features extracted at the "Conv4-4" layer of VGG19 to penalize the contrast similarity between $\hat{\mathbf{x}}$ and $\mathbf{x}$, as adopted in [39].
@@
Thus, after obtaining the trained model, in the inference stage, given a decoded image $\tilde{\mathbf{x}}$ by using Eq. (7), we can derive its enhanced image $\hat{\mathbf{x}}$ by $\hat{\mathbf{x}}=f_{\mathbf{\Theta}}^{-1}(\tilde{\mathbf{x}},\vec{\mathbf{z}})$, where $\vec{\mathbf{z}}$ refers to the latent variables sampled from Gaussian distribution $\vec{\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I})$. It can be observed that there is no additional bitrates acquirement for the enhancement processing of the decoded image.
4 Experiments
4.1 Experimental Settings
@@
Datasets. To evaluate the effectiveness of our INN-RSIC, we conducted experiments using two datasets: DOTA [40] and UC-Merced (UC-M) [41], as adopted in [18, 9]. In our experiment, the training dataset of DOTA and 80% of the UC-M training set are used for model training. During training, the images in the training dataset are randomly cropped into $128\times 128$ resolution, and several augmentation strategies, such as random horizontal flip, random vertical flip, and random crop, are adopted. To evaluate the performance of the proposed model, we randomly selected 100 images from the DOTA testing dataset, further cropped each image to a resolution of $256\times 256$, and selected 10% images from UC-M as the testing set.
@@
Implementation details. In this experiment, ELIC [15] is employed to compress the images of the training dataset. Specifically, five models are trained with parameters $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$, and these models are used to separately generate the decoded images, which are then utilized as labeled images to constrain the forward latent representation. As different compression levels correspond to various compression distortion distributions of ELIC, the labeled images under each $\lambda$ are used to train an INN-RSIC model. Besides, to train the proposed model, the widely-used AdamW [42] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 is employed for parameter optimization.  The hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are experimentally set to $1.0$, $200.0$, $0.05$, and $0.01$, respectively. The initial learning rate $\alpha$ is set to 1e-4 and the total number of epochs is set to 300, with the learning rate halved every 60 epochs until it reaches a value smaller than 1e-6. PSNR, MS-SSIM, FID [43], and LPIPS are employed as the evaluation metrics.
4.2 Performance evaluation
In this section, we conduct a comprehensive comparison between our INN-RSIC and various traditional and learning-based image compression algorithms, using the testing set of DOTA and UC-M. The benchmark includes the state-of-the-art image compression algorithms, including traditional standards such as JPEG2000 [6], BPG [7], VVC (YUV4:4:4) [8], as well as the competitive learning-based image compression algorithms like Entrorformer [21], STF [20], and ELIC [15].
The FID and LIPIS comparisons of these algorithms on the testing set of DOTA and UC-M are shown in Figure 4. It can be observed that our INN-RSIC outperforms all other methods in terms of FID and LIPIS, including traditional and deep learning-based methods, on both datasets. Furthermore, we note that at very low bpp, the improvement in both FID and LIPIS is small, which may be because the distortion of the decoded image is too high for our INN-RSIC to stably recover more details from the Gaussian distribution, which is conditioned on the decoded image. As the bpp increases, our INN-RSIC exhibits significantly superior performance in terms of FID and LPIPS.
@@ 4.2 Performance evaluation
Figure 4: Performance comparison of several image compression algorithms on the testing set of DOTA and UC-M in terms of FID and LIPIS.
(a) The testing images "P0121" and "P0216" of DOTA
(b) The testing images "baseballdiamond95" and "tenniscourt90" of UC-M
@@
Besides, Figure 5 shows the decoding results of these algorithms on the testing images of DOTA and UC-M at low and high bitrates. In Figure 5 (a), it’s noticeable that the image "P0121" decoded by the state-of-the-art traditional compression algorithm VVC exhibits obvious distortion, even with an additional 23.66% increase of bpp compared to our INN-RSIC. Moreover, the decoded image from the competitive learning-based compression algorithm STF still suffers from significant distortion, despite a 70.73% increase of bpp. At high bitrates, the decoded image "P0216" by VVC presents blurry texture details despite the higher bpp. Particularly, it is evident that the image perception quality of the decoded image by STF remains inferior to that of our INN-RSIC, even with a 65.63% increase in additional bpp. Similarly, as depicted in Figure 5 (b), the results further indicate that our INN-RSIC demonstrates impressive perception quality compared to traditional and deep learning-based image compression algorithms at both low and high bitrates.
@@
In summary, we can safely demonstrate that our INN-RSIC effectively and impressively contributes to the detailed recovery of decoded images without increasing additional bitrates.
4.3 Ablation studies
@@
Effectiveness of QM. The QM considers the impact of format mismatch on the reconstruction performance of the inverse mapping. Table 1 reports the quality metrics of the proposed model with and without QM for enhancing the decoded images of ELIC on the testing set of DOTA and UC-M. From the results, it can be seen that the proposed model with QM presents better perception quality results with comparable PSNR and MS-SSIM metrics on both datasets. The reason lies in the fact that the input images fed into our INN-RSIC are 3-channel 8-bit RGB images during the inference stage, and QM can better contribute to the proposed model to match the data type of the input images. Moreover, the lower PSNR and MS-SSIM values of the enhanced images originate from the inherent reversibility of INN. Since the ground truth is not available at the decoding end, i.e., the synthetic image is not available, we therefore use the decoded image as the inverse input to our INN-RSIC in the inference stage.

### v2

@@ III-A1 Invertible Block
Besides, as illustrated in Fig. 3(b), the forward network outputs the synthetic image $\mathbf{s}$ and the latent variable $\mathbf{z}$. Following the studies [31, 27], we assume that $\mathbf{z}$ follows a Gaussian distribution, which can be randomly sampled from the same distribution for the inverse mapping. Consequently, in the training stage, the synthetic image $\mathbf{s}$ and $\mathbf{z}$ can be merged and fed into the inverse network for model optimization. In the inference stage, the decoded image $\tilde{\mathbf{x}}$ and resampled $\vec{\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I})$ can be combined and fed into the inverse network for image enhancement.
III-A2 Channel Expansion and Haar Transformation
Within each invertible block, the channels of both its input and output signals have the same number. However, in addition to generating the synthetic image, we also need to produce an expanded output to derive the Gaussian latent variable $\mathbf{z}$. To achieve this, as shown in Fig. 3(b), we double the number of channels of the input image by using $1\times 1$ convolutional layers. In addition, Haar transformation [34] is used here to separate the compression distortion information by splitting the high and low frequencies, which are then fed into the invertible blocks.
@@ III-A2 Channel Expansion and Haar Transformation
$\mathbf{p}_{i}$ $=\operatorname{Split}(H_{D}^{-1}(\mathbf{F})),~{}~{}i\in\{1,2\},$ (4)
[…]
@@
III-A3 Conditional Generation Module (CGM)
As illustrated above, we aim to encode the compression distortion information of decoded RS images into a set of Gaussian-distributed latent variables, where the mean and variance are conditioned on the synthetic image. This conditioning enables the reconstruction process to be image-adaptive. To achieve this, as illustrated in Fig. 3(b), we introduce CGM at the end of the forward network. Specifically, to establish the dependency between $\mathbf{s}$ and $\mathbf{z}$, as depicted in Fig. 4, the output six-channel tensor $\mathbf{g}$ from the last invertible block is further divided by CGM into two parts: a three-channel synthetic image $\mathbf{s}$ and a three-channel latent variable $\tilde{\mathbf{z}}$ representing the compression distortion information.
@@ III-A3 Conditional Generation Module (CGM)
$\bm{\mu}$ $=\psi_{1}(\mathbf{s}),$ (5)
$\bm{\sigma}$ $=\exp(\psi_{2}(\mathbf{s})),$
@@
III-A4 Quantization Module (QM)
To ensure compatibility with common decoded image storage formats such as RGB (8 bits for each R, G, and B color channels), as shown in Fig. 3(b), we integrate QM after CGM. This module converts the floating-point values of the produced synthetic images into 8-bit unsigned integers by a rounding operation for quantization. However, it’s essential to acknowledge a significant obstacle: the quantization module is inherently non-differentiable. To cope with the challenge, we employ the pass-through estimator technique used in [19] to ensure that INN-RSIC is efficiently optimized during the training process. Subsequently, in the inference stage, the decoded image can be reasonably fed into the inverse network for image enhancement.
III-B Optimization Strategy
III-B1 Compression Optimization Loss
As we aim to capture the compression distortion of ELIC [19], here we provide a brief overview of its loss function. Specifically, to balance the compression ratio and the image quality of the decoded image of ELIC, the loss function can be formulated as
@@ III-B1 Compression Optimization Loss
III-B2 Forward Presentation Loss
@@
Decoded image loss
To generate the guidance images for training the forward network, ELIC is adopted to generate the labeled image $\tilde{\mathbf{x}}$, which can be formulated as
@@
Case-agnostic distribution loss
To regularize the distribution of $\mathbf{z}$, we maximize the log-likelihood of $p(\mathbf{z})$, inspired by [30]. Thus, the loss function to constrain the latent variables $\mathbf{z}$ can be formulated as
$L_{\text{case}}$ $=-\log(p(\mathbf{z})$ ) ) UNKNOWN (9)
$=-\log\left(\frac{1}{(2\pi)^{M/2}}\exp\left(-\frac{1}{2}||\mathbf{z}||^{2}\right)\right),$
where $M$ is the dimensionality of $\mathbf{z}$. This loss function penalizes the normalized latent variables $\mathbf{z}$ to follow standard Gaussian distribution. Consequently, in the inverse processing, we can randomly sample a set of Gaussian-distributed variables along with the synthetic image to derive the enhanced image.
III-B3 Reverse reconstruction loss
@@
Enhancing reconstruction loss
In theory, the synthetic image can be perfectly restored to the corresponding ground truth version through the inverse network of INN because there is no information omission. However, in practice, the decoded image is not generated by the forward network of the proposed INN-RSIC but with an image compression method. That is, the synthetic image should be stored in 8-bit unsigned integer format so that the decoded image can be used directly instead of the synthetic image here for image detail enhancement during the inference stage. To address this, we adopt QM at the end of the forward network during training. Consequently, by penalizing the discrepancy between the reconstructed image and the ground truth, we can derive the enhancing reconstruction loss given as
@@
Quality perception loss
To improve the perceptual performance of the network by estimating distances in predefined feature space rather than image space, a perceptual loss function called $L_{\text{per}}$ is used here. In other words, the feature space distance is represented by an optimization function, which serves as a driver for the network to perform image reconstruction while retaining a feature representation similar to the ground truth. Concretely, the learning perceptual image patch similarity (LPIPS) [45] is utilized to indicate the perceptual loss, which is defined as
@@
where $\operatorname{Vgg}(\cdot)$ denotes a feature function that leverages the features extracted at the “Conv4-4” layer of VGG19 to penalize the contrast similarity between $\hat{\mathbf{x}}$ and $\mathbf{x}$, as adopted in [46].
@@
Thus, after obtaining the trained model, in the inference stage, given a decoded image $\tilde{\mathbf{x}}$ by using Eq. (7), we can derive its enhanced image $\hat{\mathbf{x}}$ by $\hat{\mathbf{x}}=f_{\mathbf{\Theta}}^{-1}(\tilde{\mathbf{x}},\vec{\mathbf{z}})$, where $\vec{\mathbf{z}}$ refers to the latent variables sampled from Gaussian distribution $\vec{\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I})$. It can be observed that there is no additional bitrates acquirement for the enhancement processing of the decoded image. The training and inference stages of the proposed INN-RSIC are summarized in Algorithm 1.
1Training Stage:
2      Input: $\mathbf{x},\lambda$
9      Output: $\tilde{\mathbf{x}},\mathbf{\Theta}$
1:  Procedure: Compressor$(\mathbf{x},\lambda)$
2:     $\tilde{\mathbf{x}}=\operatorname{ELIC}(\mathbf{x},\lambda),$
3:  Return: $\tilde{\mathbf{x}}$
4:  Procedure: Enhancer$(\mathbf{x})$
5:     Initialize parameters $\mathbf{\Theta}$ of INN-RSIC with Xavier.
6:     for $\text{epoch}\leftarrow 1$ to num_epochs  do
37:        // Compute forward loss:
[…]
$\small{L_{\text{case}}=-\log\left(\frac{1}{(2\pi)^{M/2}}\exp\left(-\frac{1}{2}||\mathbf{z}||^{2}\right)\right),}$
58:        // Compute backward loss:
[…]
$\small{L_{\text{per}}=||\operatorname{Vgg}(\hat{\mathbf{x}})-\operatorname{Vgg}(\mathbf{x})||^{2},}$
79:        // Compute total loss:
$\small{L_{\text{total}}=\lambda_{1}L_{\text{dec}}+\lambda_{2}L_{\text{case}}+\lambda_{3}L_{\text{en}}+\lambda_{4}L_{\text{per}},}$
810:        //  Update $\mathbf{\Theta}$ using gradient descent:
$\small{\mathbf{\Theta}\leftarrow\mathbf{\Theta}-\alpha\nabla L_{\text{total}}(\mathbf{\Theta})},$
11:     end for
12:  Return: $\mathbf{\Theta}$
10Inference Stage:
11      Input: $\mathbf{x},\lambda$
     Output: $\tilde{\mathbf{x}},\hat{\mathbf{x}}$
1:  Procedure: Compressor$(\mathbf{x},\lambda)$
2:     $\tilde{\mathbf{x}}=\operatorname{ELIC}(\mathbf{x},\lambda),$
3:  Return: $\tilde{\mathbf{x}}$
4:  Procedure: Enhancer$(\tilde{\mathbf{x}})$
125:     Loading the trained $\mathbf{\Theta}$ for INN-RSIC.
136:     // Randomly sample $\vec{\mathbf{z}}$ and conduct inverse mapping:
[…]
7:  Return: $\hat{\mathbf{x}}$
Algorithm 1 Processing of INN-RSIC
IV Experiments
IV-A Experimental Settings
IV-A1 Datasets
@@
In this experiment, two datasets, including DOTA [47] and UC-Merced (UC-M) [48], are used for performance evaluation. Concretely, we use the training dataset of DOTA and 80% of the UC-M training set for model training. Each image is randomly cropped into the resolution of $128\times 128$. Random horizontal flip, random vertical flip, and random crop are applied for data augmentation. We randomly choose 100 images and 10% images from the DOTA testing dataset and UC-M, respectively. Then each image is centrally cropped into the resolution of $256\times 256$ as the testing set.
IV-A2 Implementation Details
@@
In our experiment, we utilize ELIC [19] to compress the images of the training dataset. Concretely, five ELIC models with parameters $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$ are used to separately derive the decoded images, which are then utilized as labeled images to constrain the forward latent representation. As different compression rates result in different compression distortion distributions of ELIC, the labeled images under each $\lambda$ are used to train the INN-RSIC model. Additionally, to train the proposed model, the widely-used AdamW [49] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 is employed for parameter optimization.  The hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are experimentally set to $1.0$, $200.0$, $0.05$, and $0.01$, respectively. We set the initial learning rate $\alpha$ to 1e-4 and the total number of epochs to 300, with the learning rate halved every 60 epochs until it reaches a value smaller than 1e-6. PSNR, MS-SSIM, FID [50], and LPIPS are adopted as the evaluation metrics.
@@
Figure 5: Performance comparison on the testing set of DOTA and UC-M. It is observed that INN-RSIC achieves a better trade-off between fidelity (i.e., higher PSNR) and perceptual quality compared to HiFiC.
IV-B Performance Evaluation
In this section, we conduct a comprehensive comparison of the proposed INN-RSIC and several competitive image compression algorithms, using the testing set of DOTA and UC-M. The benchmark includes the state-of-the-art image compression algorithms, including traditional standards such as JPEG2000 [6], BPG [7], VVC (YUV 444) [8], as well as the competitive learning-based image compression algorithms like HiFiC [29], InvComp [27], STF [24], Entrorformer [25], ELIC [19], TCM [16], and MLIC+ [17].
The performance comparisons of these algorithms on the testing set of DOTA and UC-M are shown in Fig. 5. The results indicate that the proposed INN-RSIC outperforms all other methods in terms of FID and LPIPS, with the exception of HiFiC. Although HiFiC demonstrates comparable perceptual performance, it suffers from poor image fidelity, as evidenced by its low PSNR and MS-SSIM values. Additionally, at very low bpp, the improvements in both FID and LPIPS are minimal. This may be attributed to the high distortion of the decoded image, which makes it challenging for the proposed INN-RSIC to reliably recover more details from the Gaussian distribution conditioned on the decoded image. As the bpp increases, the proposed INN-RSIC shows significantly superior performance in terms of FID and LPIPS.
(a) Testing images “P0121” and “P0216” of DOTA
(b) Testing images “baseballdiamond95” and “tenniscourt90” of UC-M
@@ IV-B Performance Evaluation
Additionally, Fig. 6 shows the decoding results of these algorithms on the testing images of DOTA and UC-M at low and high bitrates. In Fig. 6(a), it’s noticeable that the image “P0121” decoded by the state-of-the-art traditional compression algorithm VVC exhibits obvious distortion, even with an additional 23.66% increase of bpp compared to the proposed INN-RSIC. Moreover, the decoded image from the competitive learning-based compression algorithm STF still suffers from significant distortion, despite a 70.73% increase of bpp. At high bitrates, the decoded image “P0216” by VVC presents blurry texture details despite the higher bpp. Particularly, it is evident that the perception quality of the decoded images by STF remains inferior to that of the proposed INN-RSIC, even with a 65.63% increase in additional bpp. Similarly, as depicted in Fig. 6(b), the results further show that the proposed INN-RSIC demonstrates impressive perception quality compared to traditional and deep learning-based image compression algorithms at both low and high bitrates. It should be noted that, as shown in Fig. 5, since the compression performances of ELIC, TCM, and MLIC+ are comparable, for simplicity, we do not visualize the results of TCM and MLIC+.
@@
In summary, we can safely demonstrate that the proposed INN-RSIC effectively and impressively contributes to the detailed recovery of decoded images without increasing additional bitrates.
TABLE I: Quantitative comparison of the proposed INN-RSIC with and without QM or CGM
Datasets Methods FID $\downarrow$ LPIPS $\downarrow$
DOTA w/o QM 236.8280 0.3521
w/o CGM 235.6634 0.3496
INN-RSIC 195.6175 0.3218
UC-M w/o QM 174.5280 0.3600
w/o CGM 173.9155 0.3556
INN-RSIC 170.9994 0.3547
IV-C Ablation Evaluation
IV-C1 Effectiveness of QM
@@
QM considers the impact of format mismatch on the reconstruction performance of the inverse mapping. Table I reports the quality metrics of the proposed model with and without QM for enhancing the decoded images of ELIC on the testing set of DOTA and UC-M. From the results, it can be seen that the proposed model with QM presents better perception quality results with comparable PSNR and MS-SSIM metrics on both datasets. The reason lies in the fact that the input images fed into the proposed INN-RSIC are 3-channel 8-bit RGB images during the inference stage, and QM can better contribute to the proposed model to match the data type of the input images. Moreover, the lower PSNR and MS-SSIM values of the enhanced images originate from the inherent reversibility of INN. Since the ground truth is not available at the decoding end, i.e., the synthetic image is not available, we therefore use the decoded image as the inverse input to the proposed INN-RSIC in the inference stage.
IV-C2 Effectiveness of CGM